# KNighter: Transforming Static Analysis with LLM-Synthesized Checkers

### v2

@@ Abstract.
We present KNighter, the first approach that unlocks scalable LLM-based static analysis by automatically synthesizing static analyzers from historical bug patterns. Rather than using LLMs to directly analyze massive systems, our key insight is leveraging LLMs to generate specialized static analyzers guided by historical patch knowledge. KNighter implements this vision through a multi-stage synthesis pipeline that validates checker correctness against original patches and employs an automated refinement process to iteratively reduce false positives. Our evaluation on the Linux kernel demonstrates that KNighter generates high-precision checkers capable of detecting diverse bug patterns overlooked by existing human-written analyzers. To date, KNighter-synthesized checkers have discovered 92 new, critical, long-latent bugs (average 4.3 years) in the Linux kernel; 77 are confirmed, 57 fixed, and 16 have been assigned CVE numbers. This work establishes an entirely new paradigm for scalable, reliable, and traceable LLM-based static analysis for real-world systems via checker synthesis.
@@ 1. Introduction
We implement our approach in a tool, KNighter, the first fully automated pipeline for synthesizing static analyzers, built upon the open-source Clang Static Analyzer (CSA) (clang-static-analysis, ). Although the methodology generalizes to different systems, our evaluation targets the Linux kernel, one of the most fundamental software systems. From 61 historical patch commits, KNighter successfully synthesizes 37 high-quality checkers with a false positive rate of about 35%. Furthermore, the vulnerabilities detected are orthogonal to those found by existing expert-written analyzers (smatch, ). Demonstrating practical impact, our synthesized checkers have uncovered 92 new, long-latent vulnerabilities (average 4.3 years) in the Linux kernel, resulting in 77 developer confirmations, 57 fixes, and 16 CVEs. These findings validate our approach’s efficacy and its contribution to system reliability.
@@
  • Evaluation. We evaluate KNighter by synthesizing checkers for the patch commits spanning multiple bug categories. Our results demonstrate that KNighter generates high-precision checkers with practical false positive rates.
  • Real-world impact. KNighter-generated checkers have discovered 92 new, long-latent (average 4.3 years) bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 assigned CVE numbers—demonstrating its practical impact on system reliability and security.
@@ 3. Design
Overview. KNighter leverages agentic workflow to process patch commits for static analyzer synthesis, as illustrated in Figure 3. It operates in two phases: checker synthesis (§ 3.1) and checker refinement (§ 3.2). In the checker synthesis phase, KNighter analyzes the input patch to identify bug patterns (§ 3.1.1), synthesizes a detection plan (§ 3.1.2), and implements a checker using CSA (§ 3.1.3). If compilation errors occur, a syntax-repair agent automatically repairs them based on the error messages. This phase concludes with the generation of valid checkers (§ 3.1.4). In the subsequent checker refinement phase, these valid checkers are deployed to scan the entire codebase for potential bugs. When bug reports are generated, a bug-triage agent evaluates them for false positives, and KNighter refines the checker accordingly. If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and the filtered reports as potential bugs for review.
@@ 3.1.2. Plan Synthesis
(a) Prompt template for bug pattern analysis
(b) Prompt template for plan synthesis.
Figure 5. Prompt templates used by KNighter.
3.1.3. Implementation and Syntax Repair
@@ 3.2. Checker Refinement
Our refinement pipeline addresses these challenges methodically. First, to manage report complexity, we distill generated bug reports to their essential components—primarily the “relevant lines” highlighted by the static analyzer (e.g., CSA (clang-static-analysis, )) and the corresponding trace path—stripping extraneous context while preserving critical diagnostics. Second, to navigate the complexity of analysis and modification, we employ specialized LLM-based agents. A triage agent classifies each distilled report, focusing strictly on pattern alignment (i.e., does the report match the original input bug’s pattern?) rather than general code correctness.
@@ 5.1.1. Checker Synthesis
KNighter 92 77 57 15 16
(a) Bug summary statistics detected by KNighter.
(b) Number of bugs in each type.
(c) Number of bugs in each subsystem.
(d) Number of bugs with different lifetimes.
Figure 9. The details of newly detected bugs by KNighter. Subfigure (9(a)) summarizes overall status, while (9(b)), (9(c)), and (9(d)) show breakdowns by type, subsystem, and lifetime.
@@ 5.2. RQ2: Detected Bugs
To date, static analyzers synthesized by KNighter have identified 92 new bugs in the Linux kernel. As summarized in Figure 9(a), developer confirmation has been received for 77 of these bugs. Among those confirmed, 57 have already been fixed. The remaining 15 bugs are currently awaiting developer review (calculated as Total - Confirmed). Notably, 16 of the discovered bugs have been assigned CVE numbers, showing their practical security impact.
@@
The checkers for bug detection originate from two sources: (i) the initial 61 manually collected commits used for evaluation in § 3.1 across diverse bug types (shown in Table 1), and (ii) an additional set of 100 commits automatically collected using keywords related to Null-Pointer-Dereference to further explore that specific bug class. Figure 9(b) and Figure 9(c) show the bug distribution from these sources—with the light blue for bugs from the initial evaluation set and the light purple for those from the auto-collected set.
Bug types. Analysis of checkers from the manually collected commits reveals that KNighter can find a wide range of bug types (in Figure 9(b)). Among these, Null-Pointer-Dereference bugs are the most prevalent, showing the KNighter’s strength in this area. In response, we expanded our effort by automatically collecting commits related to Null-Pointer-Dereference, which yielded an additional 30 bugs.
Bug location. As shown in Figure 9(c), the detected bugs span various Linux kernel subsystems. The majority appear in the drivers subsystem (62 out of 92), reflecting its large footprint in the kernel (bursey2024syzretrospector, ). Additionally, 14 and 7 bugs were identified in the sound and net subsystems, respectively. Notably, 2 bugs were found in the samples directory—an area that provides example usage for kernel developers and where correctness is especially critical.
Bug lifetime. Figure 9(d) illustrates the distribution of bug lifetimes. Notably, the average bug lifetime is 4.3 years, and 26 bugs had existed for over five years before detection. This indicates that the vulnerabilities uncovered by KNighter are difficult to detect and remain latent for extended periods, underscoring the effectiveness of our approach.
5.2.2. Case Study.
Here are examples of vulnerabilities detected by KNighter.
@@ 6.1. Traditional Static Analysis
Our synthesis approach. In contrast to the above, our work, KNighter, uses LLMs to synthesize the static analyzer itself directly from historical patch information. Instead of relying on predefined rules or inferring specifications for existing checkers, KNighter learns both the bug pattern and the corresponding detection logic, generating a new checker automatically. This approach fundamentally differs by automating the analyzer creation process, potentially offering greater adaptability to new bug types compared to traditional methods or specification inference. This distinction may also make KNighter more versatile; while specification-inference tools excel at finding bugs related to the interfaces they model, KNighter can synthesize checkers for a broader range of issues demonstrated in patches, including problems not strictly tied to interface misuse, such as certain integer overflows or complex logic errors. Furthermore, although KNighter currently generates CSA checkers, the core idea of LLM-based checker synthesis is potentially generalizable to target other static analysis frameworks, given the broad training data of modern LLMs. Thus, KNighter complements existing approaches by providing a mechanism to automatically generate analyzers for newly observed bug patterns found in patches.
@@ 6.2. LLM-Based Static Analysis
LLM-augmented static analysis. Some techniques use LLMs to assist existing tools by automating tasks previously requiring manual effort. For example, LLMs can infer taint specifications for external APIs (IRIS (li2024llm, )), generate post-constraints to prune analysis paths (LLift (li2024enhancing, )), or infer resource-handling intentions (InferROI (wang2024resourceleak, )). While helpful, these methods still fundamentally depend on a substantial, human-developed analyzer core. This can limit their ability to easily adapt or generalize to detect novel bug types beyond those originally targeted by the core analyzer’s design.
@@
Our synthesis approach. In contrast, KNighter uses LLMs to synthesize the entire static analyzer from patches. Unlike LLM-augmented tools, this minimizes reliance on pre-existing manual analyzer development, enhancing adaptability for diverse bug types. Unlike direct LLM analysis, KNighter generates efficient, reusable static checkers, avoiding the high costs and scalability issues of scanning massive systems directly with LLMs. This makes KNighter a practical approach for applying LLM intelligence to large-scale static analysis, complementing other methods by automatically creating analyzers for new bug patterns. Checkers generated by KNighter could also potentially benefit from integration with LLM-augmented techniques.
@@ 7. Conclusion
This paper introduces KNighter, a novel approach that transforms how LLMs can contribute to static analysis for complex systems like the Linux kernel. By synthesizing static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive systems. KNighter’s practical impact is shown by the discovery of 92 new, long-latent bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 16 CVE assigned.

### v3

@@ Abstract.
We present KNighter, the first approach that unlocks scalable LLM-based static analysis by automatically synthesizing static analyzers from historical bug patterns. Rather than using LLMs to directly analyze massive systems, our key insight is leveraging LLMs to generate specialized static analyzers guided by historical patch knowledge. KNighter implements this vision through a multi-stage synthesis pipeline that validates checker correctness against original patches and employs an automated refinement process to iteratively reduce false positives. Our evaluation on the Linux kernel demonstrates that KNighter generates high-precision checkers capable of detecting diverse bug patterns overlooked by existing human-written analyzers. To date, KNighter-synthesized checkers have discovered 92 new, critical, long-latent bugs (average 4.3 years) in the Linux kernel; 77 are confirmed, 57 fixed, and 30 have been assigned CVE numbers. This work establishes an entirely new paradigm for scalable, reliable, and traceable LLM-based static analysis for real-world systems via checker synthesis.
Static Analysis, Large Language Models
journalyear: 2025copyright: ccconference: ACM SIGOPS 31st Symposium on Operating Systems Principles; October 13–16, 2025; Seoul, Republic of Koreabooktitle: ACM SIGOPS 31st Symposium on Operating Systems Principles (SOSP ’25), October 13–16, 2025, Seoul, Republic of Koreadoi: 10.1145/3731569.3764827isbn: 979-8-4007-1870-0/2025/10ccs: Security and privacy Systems securityccs: Software and its engineering Automated static analysis
@@ 1. Introduction
We implement our approach in a tool, KNighter, the first fully automated pipeline for synthesizing static analyzers, built upon the open-source Clang Static Analyzer (CSA) (clang-static-analysis, ). While the methodology generalizes to different systems, we target the Linux kernel, one of the most fundamental software systems. In the evaluation of 61 diverse bug-fix patches, KNighter synthesized the high-quality checkers for 61% of them, achieving a false positive rate of about 35% aided by the triage agent. Demonstrating practical impact, KNighter has uncovered 92 new, long-latent vulnerabilities (average 4.3 years) in the Linux kernel, resulting in 77 developer confirmations, 57 fixes, and 30 CVEs. Furthermore, the vulnerabilities detected are orthogonal to those found by existing expert-written analyzers (smatch, ). These findings validate our approach’s efficacy and its contribution to system reliability.
@@
  • Evaluation. We demonstrate that KNighter successfully synthesizes effective checkers from the Linux kernel bug-fix patches across various bug categories, achieving practical false positive rates.
  • Real-world impact. KNighter-generated checkers have discovered 92 new, long-latent (average 4.3 years) bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 30 assigned CVE numbers—demonstrating its practical impact on system reliability and security.
@@ 3. Design
Overview. KNighter leverages agentic workflow to process patch commits for static analyzer synthesis, as illustrated in Figure 3. It operates in two phases: checker synthesis (§ 3.1) and checker refinement (§ 3.2). In the checker synthesis phase, KNighter analyzes the input patch to identify bug patterns (§ 3.1.1), synthesizes a detection plan (§ 3.1.2), and implements a checker using CSA (§ 3.1.3). If compilation errors occur, a syntax-repair agent automatically repairs them based on the error messages. This phase concludes with the generation of valid checkers (§ 3.1.4). In the subsequent checker refinement phase, these valid checkers are deployed to scan the entire codebase for potential bugs. When bug reports are generated, a triage agent evaluates them for false positives, and KNighter refines the checker accordingly. If the scan produces a manageable number of reports with a low false positive rate, KNighter presents the plausible checkers and their filtered reports as potential bugs for review.
@@ 22        Continue
(a) Prompt template for bug pattern analysis
(b) Prompt template for plan synthesis.
(c) Prompt template for report triage.
Figure 5. Simplified prompt templates used by KNighter.
@@ 3.1.2. Plan Synthesis
3.1.3. Analyzer Implementation and Syntax Repair
Figure 6. Pre-defined checker template for CSA.
@@ 3.2. Checker Refinement
Our refinement pipeline addresses these challenges methodically. First, to manage report complexity, we distill generated bug reports to their essential components—primarily the “relevant lines” highlighted by the static analyzer (e.g., CSA (clang-static-analysis, )) and the corresponding trace path—stripping extraneous context while preserving critical diagnostics. Second, to navigate the complexity of analysis and modification, we employ specialized LLM-based agents. A triage agent classifies each distilled report, focusing strictly on alignment with the target bug pattern (rather than general code correctness); the prompt template is shown in Figure 5(c).
@@ 5.1.1. Checker Synthesis
KNighter 92 77 57 15 30
(a) Number of bugs in each type.
(b) Number of bugs in each subsystem.
(c) Number of bugs with different lifetimes.
(d) Number of bugs detected by each commit.
Figure 9. Details of new bugs. Subfigures (9(a)), (9(b)), (9(c)), and (9(d)) show breakdowns by type, subsystem, lifetime, and commit.
@@ 5.2. RQ2: Detected Bugs
To date, static analyzers synthesized by KNighter have identified 92 new bugs in the Linux kernel. As summarized in Table 2, developer confirmation has been received for 77 of these bugs. Among those confirmed, 57 have already been fixed. The remaining 15 bugs are currently awaiting developer review (calculated as Total - Confirmed). Notably, 30 of the discovered bugs have been assigned CVE numbers, showing their practical security impact.
The checkers for bug detection originate from two sources: (i) the initial 61 manually collected commits used for evaluation in § 3.1 across diverse bug types (shown in Table 1), and (ii) an additional set of 100 commits automatically collected using keywords related to Null-Pointer-Dereference to further explore that specific bug class. Figure 9(a) and Figure 9(b) show the bug distribution from these sources—with the light blue for bugs from the initial evaluation set and the light purple for those from the auto-collected set.
Bug types. Analysis of checkers from the manually collected commits reveals that KNighter can detect a diverse range of bug types (see Figure 9(a)). Null-Pointer-Dereference bugs are the most prevalent, highlighting KNighter ’s strength in this area. In response, we expanded our effort by automatically collecting commits related to Null-Pointer-Dereference, which yielded an additional 30 bugs.
Bug location. As shown in Figure 9(b), the detected bugs span various Linux kernel subsystems. The majority appear in the drivers subsystem (67 out of 92), reflecting its large footprint in the kernel (bursey2024syzretrospector, ). Additionally, 10 and 7 bugs were identified in the sound and net subsystems, respectively. Notably, 2 bugs were found in the samples directory—an area that provides example usage for kernel developers and where correctness is especially critical.
Bug lifetime. Figure 9(c) illustrates the distribution of bug lifetimes. Notably, the average bug lifetime is 4.3 years, and 26 bugs had existed for over five years before detection. This indicates that the vulnerabilities uncovered by KNighter are difficult to detect and remain latent for extended periods, underscoring the effectiveness of our approach.
Bug distribution over commits. Synthesized checkers from 39 bug-fix commits uncovered new bugs (2.4 each on average), with a long-tail skew (as shown in Figure 9(d)): five checkers each found five or more bugs. In general, checkers derived from recurring error patterns yield higher counts, while those from specialized fixes surface fewer, yet still valuable, findings. This suggests that KNighter can learn and propagate impactful patterns beyond their original contexts, producing a mix of broad-coverage and high-yield checkers.
5.2.2. Case Study.
Here are examples of vulnerabilities detected by KNighter.
@@ 5.4.1. Bug Triage Agent
We also evaluated 5-way self-consistency via majority voting (wang2022self, ). For each report, we ran our triage agent five independent times and labeled the report as a “bug” only if the agent made that prediction in at least $t$ of the runs. Compared to our single-sample baseline, which identified 7 TPs and 22 FPs, majority voting did not offer a significant improvement. Using a threshold of $t=3$ kept the TP count at 7 but slightly increased FPs to 24. A stricter threshold of $t=4$ also resulted in 7 TPs while reducing FPs to 20. Ultimately, majority voting only slightly shifted the false positive count in this setting. There could be two potential reasons. For TPs, the agent likely identifies these bugs with high confidence, meaning a single run is sufficient to detect all of them. For FPs, the agent’s predictions are likely less confident and more inconsistent across different runs.
@@ 5.4.2. Ablation Study
6. Limitations and Discussion
Limitations. KNighter faces challenges with highly complex bug patterns, particularly those involving state-machine reasoning, such as use-after-free, and concurrency issues that require analyzing multi-threaded code and locking schemes. Synthesizing precise checkers for these issues is difficult for today’s LLMs, as it requires a sophisticated understanding of a program’s temporal and interprocedural behavior. To address this, a promising future direction is to improve the analysis agent to automatically abstract complex bug reports into more formal state-machine representations. This would enable the synthesis of precise, state-aware checkers that can trace the specific sequences of operations leading to a bug. A second, complementary strategy is to focus on selectively identifying high-quality, canonical bug fixes. This involves developing methods to filter out patches that are overly complex or specific, allowing KNighter to learn from reusable, idiomatic repair patterns and improve the overall performance of checker synthesis.
Generability. KNighter employs a general three-stage workflow: LLM-driven synthesis of checkers from bug-fix patches, patch-grounded validation, and triage/refinement to reduce false positives. While the Linux kernel served as an ideal and challenging initial target due to its scale, complexity, and importance, our approach is not fundamentally tied to it. The workflow’s applicability extends to any project that meets two criteria: a version history with bug-fix commits for learning, and a static analysis framework to serve as a synthesis target (e.g., Chromium (chromium, )). Furthermore, the implementation is not limited to C/C++. Although we demonstrated its use with the Clang Static Analyzer (clang-static-analysis, ), the pipeline can be readily adapted—via a small set of few-shot examples—to generate checkers or rules for other ecosystems (e.g., CodeQL (codeql, ) or Semgrep (semgrep, )).
@@ 7. Related Work
This section mainly discusses related work on both traditional and LLM-based static analysis, contextualizing the contributions of KNighter. While our work focuses on synthesizing static analyzers that inspect source code without execution, it is worth noting that checker synthesis has also been explored for generating dynamic runtime checks (t2c, ; lou2022demystifying, ; TrainCheckOSDI2025, ). These approaches are fundamentally different from KNighter. Dynamic checkers typically learn rules from test executions (t2c, ) to detect bugs based on concrete runtime states, such as capturing semantic “grey-failures”. In contrast, KNighter learns from source code patches to create static checkers that can cover all potential execution paths, including corner cases rarely triggered during program execution.
@@ 7.1. Traditional Static Analysis
Deviation-based specification inference. Methods such as (engler2001bugs, ; min2015cross, ; ahmadi2021finding, ) infer specifications by assuming the majority of uses are correct and flagging deviations. While effective in some settings, classic systems rely on a small, fixed set of rule templates plus probabilistic clustering/ranking, which constrains bug-pattern coverage and—without strong semantic post-checks or refinement—often yields higher FPR (engler2001bugs, ; min2015cross, ). FICS (ahmadi2021finding, ) reduces template dependence via ML grouping of functionally similar code, but still rests on the majority-correct assumption and similarity heuristics, which can miscluster code and introduce noise.
@@
Our synthesis approach. In contrast to the above, KNighter uses LLMs to synthesize the static analyzer itself directly from historical patch information. Instead of relying on predefined rules or inferring specifications for existing checkers, KNighter learns both the bug pattern and the corresponding detection logic, generating a new checker automatically. This approach fundamentally differs by automating the analyzer creation process, potentially offering greater adaptability to new bug types compared to traditional methods or specification inference. This distinction may also make KNighter more versatile; while specification-inference tools excel at finding bugs related to the interfaces they model, KNighter can synthesize checkers for a broader range of issues demonstrated in patches, including problems not strictly tied to interface misuse, such as certain integer overflows or complex logic errors. Furthermore, although KNighter currently generates CSA checkers, the core idea of LLM-based checker synthesis is potentially generalizable to target other static analysis frameworks, given the broad training data of modern LLMs. Thus, KNighter complements existing approaches by providing a mechanism to automatically generate analyzers for newly observed bug patterns found in patches.
@@ 7.2. LLM-Based Static Analysis
LLM-augmented static analysis. Some techniques use LLMs to assist existing tools by automating tasks previously requiring manual effort. For example, LLMs can infer taint specifications for external APIs (IRIS (li2024llm, ), Artemis (ji2025artemis, )), generate post-constraints to prune analysis paths (LLift (li2024enhancing, )), or infer resource-handling intentions (InferROI (wang2024resourceleak, )). While helpful, these methods still fundamentally depend on a substantial, human-developed analyzer core, e.g., the manually defined rules for taint propagation (ji2025artemis, ). This can limit their ability to easily generalize to detect novel bug types beyond those originally targeted by the core analyzer’s design.
@@
Our synthesis approach. In contrast, KNighter uses LLMs to synthesize the entire static analyzer from patches. Unlike LLM-augmented tools, this minimizes reliance on pre-existing manual analyzer development, enhancing adaptability for diverse bug types. Unlike direct LLM analysis, KNighter generates efficient, reusable static checkers, avoiding the high costs and scalability issues of scanning massive codebases directly with LLMs. A very recent concurrent work, MoCQ (mocq, ), also explores checker/query synthesis, but it focuses on general bug patterns and relies on manual examples for validation. KNighter distinguishes itself not only by automatically inferring specific, nuanced patterns from patches but also by incorporating a closed-loop triage-refinement pipeline to iteratively improve checker precision. This fully automated refinement process makes our approach more scalable for detecting complex defects in system software, establishing it as a practical paradigm for applying LLM intelligence to large-scale static analysis.
@@ 8. Conclusion
This paper introduces KNighter, a novel approach that transforms how LLMs can contribute to static analysis for complex systems like the Linux kernel. By synthesizing static analyzers rather than directly analyzing code, KNighter bridges the gap between LLMs’ reasoning capabilities and the practical constraints of analyzing massive systems. KNighter’s practical impact is shown by the discovery of 92 new, long-latent bugs in the Linux kernel, with 77 confirmed, 57 fixed, and 30 CVE assigned.
@@
Acknowledgments
We are grateful to the anonymous reviewers for their valuable feedback that helped to improve this paper. This work was partially supported by NSF grant CCF-2131943.